# Total Failure of Approachability at Successors of Singulars of Countable Cofinality

### v1

@@ 4. Collapsing without Approachability
By induction. The beginning is clear. Assume the statement has been shown for $J_{0},\dots,J_{n-1}$ and let $J_{n}$ be some ideal. Let $A\in(J_{0}\otimes\dots\otimes J_{n})^{+}$. It follows that for $(J_{0}\otimes\dots\otimes J_{n-1})^{+}$ many $s\in\prod_{k<n}X_{k}$, $\{x\in X_{n}\;|\;s^{\frown}x\in A\}\in J_{n}^{+}$. By the inductive hypothesis, let $p$ be a $(J_{k})_{k<n}$-Laver-tree such that whenever $s\in p$, $\{x\in X_{n}\;|\;s^{\frown}x\in A\}\in J_{n}^{+}$. Let $p^{\prime}$ be the tree of all $n+1$-element sequences $s$ with $s\upharpoonright n\in p$ and $(s\upharpoonright n,s(n))\in A$. then $p^{\prime}$ is a $(J_{k})_{k\leq n}$-Laver-tree and $p^{\prime}]\subseteq A$. ∎
@@
where $\subseteq^{*}$ is defined with regards to $I_{0}\otimes\dots\otimes I_{n-1}$.
@@
$\{t\in\operatorname{osucc}_{q_{2}}^{n}(\emptyset)\;|\;\operatorname{osucc}_{q_{2}}(t)\not\subseteq^{*}\operatorname{osucc}_{q_{i}}(t)\}\in I_{0}\otimes\dots\otimes I_{n-1}$
@@
$\{t\in\operatorname{osucc}_{p}^{n}(\emptyset)\;|\;\operatorname{osucc}_{p}(t)\in I_{n}\}\subseteq A\in I_{0}\otimes\dots\otimes I_{n-1}$
@@
$\{r\in\operatorname{osucc}_{p}^{n}(t)\;|\;\operatorname{osucc}_{p}^{n}(t^{\frown}r)\in I_{|t|+n}\}\in I_{|t|}\otimes\dots\otimes I_{|t|+(n-1)}$
Let $q$ consist of all those $t\in q^{\prime}$ such that every initial segment of $t$ is also in $q^{\prime}$. We claim that $q\in\mathbb{L}$ and $\operatorname{stem}(q)=\emptyset$. By Claim 2 and Claim 3, $\emptyset\in q$ and so $q$ is nonempty. It is clearly closed under initial segments by construction. So we have to show that it splits wide enough. Let $t\in q$. By assumption, $\operatorname{osucc}_{p}(t)\in I_{|t|}^{+}$. Additionally, $\operatorname{osucc}_{q}(t)$ consists of all $x\in\operatorname{osucc}_{p}(t)$ such that $\operatorname{osucc}_{p}(t^{\frown}x)\in I_{|t|+1}^{+}$ and for every $n$,
$\{r\in\operatorname{osucc}_{p}^{n}(t^{\frown}x)\;|\;\operatorname{osucc}_{p}(t^{\frown}x^{\frown}r)\in I_{|t|+1+n}\}\in I_{|t|+1}\otimes\dots\otimes I_{|t|+n}$
@@
$\{r\in\operatorname{osucc}_{p}^{n}(t^{\frown}x)\;|\;\operatorname{osucc}_{p}(t^{\frown}x^{\frown}r)\in I_{|t|+1+n}\}\in(I_{|t|+1}\otimes\dots\otimes I_{|t|+n})^{+}$
@@
contradicting the fact that $t\in q$ (more precisely, the second requirement for $n=1$). In the second case, we have that the set
@@
$\{x\in[\kappa_{|t|}]^{<\mu^{+}}\;|\;\{r\in\prod_{i\in[|t|+1,|t|+n]}[\kappa_{i}]^{<\mu^{+}}\;|\;(x,r)\in A\}\in(I_{|t|}\otimes\dots\otimes I_{|t|+n})^{+}\}\in I_{|t|}^{+}$
In other words, $A\in(I_{|t|}\otimes(I_{|t|+1}\otimes\dots\otimes I_{|t|+n}))^{+}=(I_{|t|}\otimes\dots\otimes I_{|t|+n})^{+}$ (the equality follows from Lemma 4.6). But this clearly contradicts the fact that $t\in q$ (more precisely, the second requirement for $n+1$).
@@
The following is our crucial strengthening of the main crux of our earlier work (see [JL25, Theorem 4.3]). Instead of showing that any function $\dot{F}$ of the stated form has an initial segment outside of the ground model we are showing that any such function has an initial segment which is not even covered by a small set which is in the ground model. This directly implies that $\kappa^{*}$ does not become approachable without relying on our collapsing forcing not adding short functions into small sets.
@@
$p\upharpoonright(\operatorname{stem}(p)^{\frown}x_{0})\Vdash\dot{F}(\check{i})\in\check{A}_{x_{0}}$
@@
$(p_{x},\dot{q}_{x})\Vdash\dot{F}(\check{i}_{x})\in\check{A}_{x}$
@@
$A_{x}:=\{\alpha\;|\;\exists s\in p_{x}(|s|=n_{x}\wedge(p_{x}\upharpoonright s,\dot{q}_{x})\Vdash\dot{F}(\check{i}_{x})=\check{\alpha})\}$
@@
$(p_{x}^{\prime},\dot{q}_{x}^{\prime})\Vdash\dot{F}(\check{i}_{x})\notin\check{B}$
@@
Let $G\colon[p]\to\mu$ be the function mapping a branch $b\in[p]$ to $\sup_{n}(i_{b\upharpoonright n})$. For every $\alpha<\mu$, the set $\{b\in[p]\;|\;G(b)<\alpha\}$ is closed in $[p]$ (with regards to the tree topology), so by the Rubin-Shelah Theorem (see e.g. [She17, Chapter XI, Lemma 3.5 (1)]) and the fact that every $I_{n}$ is $\mu^{+}-$complete, we can find $p^{\prime}\leq_{0}p$ such that $G$ is bounded in $p_{0}$, say with value $i$. We claim that $(p^{\prime},\dot{q}^{\prime})$ and $i$ work.
@@ 5. Non-approachable Points of a Single Cofinality
Let $\kappa$ be a supercompact cardinal and $l\colon\kappa\to V_{\kappa}$ a Laver function. Let $\mu<\kappa$ be regular. We define the poset $(\mathbb{A}(\mu,\kappa),\leq,\leq_{0})$ as the limit of an Easton-support Magidor iteration $(\mathbb{P}_{\alpha},\dot{\mathbb{Q}}_{\alpha})_{\alpha<\kappa}$ with the following iterands (we let $\dot{\mathbb{Q}}_{\alpha}$ be trivial for $\alpha\leq\mu$):
@@
We fix the following notation: Given $\xi<\kappa$, let $\mathbb{A}(\xi)$ consist of all $f\in\mathbb{A}(\mu,\kappa)$ such that $\operatorname{dom}(f)\subseteq\xi$, ordered as a suborder of $\mathbb{A}(\mu,\kappa)$. It is clear that whenever $\xi<\kappa$, $\mathbb{A}(\xi)$ is a regular suborder of $\mathbb{A}$ (and restriction to $\xi$ is a projection from $\mathbb{A}$ to $\mathbb{A}(\xi)$). So, for any $\xi<\kappa$, we can let $G(\xi)$ be the $\mathbb{A}(\xi)$-generic filter induced by $G$. It is clear that $G(\xi)=G\cap\mathbb{A}(\xi)$ whenever $\xi<\kappa$ and that $\pi[G]=G(\nu)$.
We also fix the following ordinals: Let $\gamma:=\sup(M\cap\delta^{*})$ and $\delta^{*}_{-}:=\operatorname{otp}(M\cap\delta^{*})$. Then, by Lemma 1.11, $\delta^{*}_{-}=\operatorname{cf}(\gamma)$ is the successor of a singular cardinal of countable cofinality and $d$ induces a normal coloring $e$ on $\delta^{*}_{-}$ such that in any outer model, $\gamma$ is $d$-approachable if and only if $\delta^{*}_{-}$ is $e$-approachable. For $n\in\omega$, let $\delta_{n}^{-}:=\operatorname{otp}(M\cap\delta_{n})$. Then each $\delta_{n}^{-}$ is a regular cardinal and $\delta^{*}_{-}=(\sup_{n}\delta_{n}^{-})^{+}$.
We now prove that $M$ satisfies (2). It follows that $\mathbb{A}(\nu)$ is $\nu$-cc and so in $V[G(\nu)]$, $\nu$, every $\delta_{n}^{-}$ and $\delta^{*}_{-}$ are still regular cardinals. Moreover, $l(\nu)=\pi_{N}((\dot{\mathbb{P}},(\delta_{n})_{n\in\omega}),\delta^{*})=(\pi_{N}(\dot{\mathbb{P}}),(\delta_{n}^{-})_{n\in\omega},\delta^{*}_{-})$ and the following holds (since $\pi_{N}$ is an isomorphism and the considered statements are sufficiently elementary, see the discussion after Lemma 1.13):
@@
    $\delta^{*}_{-}>\nu$ is the successor of a cardinal of countable cofinality.
@@
    $(\delta_{n}^{-})_{n\in\omega}$ converges to the predecessor of $\delta^{*}_{-}$
@@
$\mathbb{A}(\nu+3)\cong\mathbb{A}(\nu)*\pi_{N}(\mathbb{P})*\mathbb{L}(\check{\mu},(\check{\delta}_{n}^{-},\dot{I}_{n},\dot{B}_{n})_{n\in\omega})*\dot{\operatorname{Coll}}(\check{\mu},\check{\delta}^{*})$
@@
We want to show that $\operatorname{cf}(\gamma)=\mu$ and $\gamma$ is not $d$-approachable in $V[G(\nu+3)]$. First of all, $e$ is still a normal subadditive coloring on $\delta^{*}_{-}$ after forcing with $\pi_{N}(\dot{\mathbb{P}})^{G\cap\mathbb{A}(\nu)}$ (since $\delta^{*}_{-}$ and every $\delta_{n}$ is preserved). By Theorem 4.12, $\operatorname{cf}(\delta^{*}_{-})=\mu$ and $\delta^{*}_{-}$ is not $e$-approachable in $V[G(\nu+3)]$. Hence, still in $V[G(\nu+3)]$, $\operatorname{cf}(\gamma)=\mu$ and $\gamma$ is not $d$-approachable.
@@
We now verify that this is preserved when going to $V[G]$. $V[G]$ is a forcing extension of $V[G(\nu+3)]$ using a Prikry-type poset $(\mathbb{A}(\kappa\smallsetminus\nu+3),\leq,\leq_{0})$ such that $(\mathbb{A}(\kappa\smallsetminus\nu+3),\leq_{0})$ is $\mu-$closed and has direct $\mu$-decidability (using the same argument as for Lemma 5.2 (2)). Since there is a strictly increasing and cofinal function from $\mu$ to $\gamma$ in $V[G(\nu+1)]$, the cofinalities of $\mu$ and $\gamma$ are equal in any outer model. Since $\mu$ is a regular cardinal in $V[G]$, $\operatorname{cf}(\gamma)=\mu$ in $V[G]$.
@@
We now prove that $\gamma$ does not become $d$-approachable in $V[G]$. In $V[G(\nu+3)]$, fix a function $\Delta\colon\mu\to\gamma$ which is increasing and cofinal. By Fact 1.5, $\gamma$ is $d$-approachable in $V[G]$ if and only if there is an unbounded subset of $\operatorname{im}(\Delta)$ on which $d$ is bounded. So assume toward a contradiction that $\dot{\Gamma}$ is an $(\mathbb{A}(\kappa\smallsetminus\nu+3),\leq)$-name for an increasing function from $\mu$ into $\operatorname{im}(\Delta)$ such that $d$ is bounded on $\operatorname{im}(\dot{\Gamma})$ (say, with value $\check{n}$), forced by some $a\in\mathbb{A}(\kappa\smallsetminus\nu+3)$. We now construct an increasing sequence $(\alpha_{i})_{i<\mu}$ of elements of $\mu$ and a $\leq_{0}$-decreasing sequence $(a_{i})_{i<\mu}$ (with $a_{0}\leq_{0}a$) of elements of $\mathbb{A}(\kappa\smallsetminus\nu+3)$ such that for any $i<\mu$, $a_{i}\Vdash\dot{\Gamma}(\check{i})=\check{\Delta}(\check{\alpha}_{i})$. If the sequences have been constructed for $j<i$, first let $a_{i}^{\prime}$ be a $\leq_{0}$-lower bound of $(a_{j})_{j<i}$ (or $a_{i}^{\prime}:=a$ if $i=0$). Then $a_{i}^{\prime}\Vdash\exists\alpha<\mu(\dot{\Gamma}(\check{i})=\check{\Delta}(\alpha))$ so by the maximum principle and the direct $\mu$-decidability of $(\mathbb{A}(\kappa\smallsetminus\nu+3),\leq,\leq_{0})$ we can find $\alpha_{i}$ and $a_{i}\leq_{0}a_{i}^{\prime}$ such that
$a_{i}\Vdash\dot{\Gamma}(\check{i})=\check{\Delta}(\check{\alpha}_{i})$
Since $\dot{\Gamma}$ is forced to be and $\Delta$ is increasing, $\alpha_{i}>\alpha_{j}$ for all $j<i$.
@@
Now we claim that $\{\Delta(\alpha_{i})\;|\;i<\mu\}$ is a cofinal subset of $\gamma$ on which $d$ is bounded with value $n$, obtaining a contradiction as that set is in $V[G(\nu+3)]$. First observe that, because the sequence $(\alpha_{i})_{i<\mu}$ is increasing, the set $\{\alpha_{i}\;|\;i<\mu\}$ is cofinal in $\mu$ and so, since $\Delta$ is increasing and cofinal, $\{\Delta(\alpha_{i})\;|\;i<\mu\}$ is cofinal in $\gamma$. Lastly, whenever $j<i<\mu$, we have
@@
so $d(\Delta(\alpha_{j}),\Delta(\alpha_{i}))\leq n$.

### v2

@@ 4. Collapsing without Approachability
We prove both directions simultaneously by induction with the beginning being clear.
Assume the statement has been shown for $J_{0},\dots,J_{n-1}$ and let $J_{n}$ be some ideal.
Let $A\in(J_{0}\otimes\dots\otimes J_{n})^{+}$. It follows that for $(J_{0}\otimes\dots\otimes J_{n-1})^{+}$ many $s\in\prod_{k<n}X_{k}$, $\{x\in X_{n}\;|\;s^{\frown}x\in A\}\in J_{n}^{+}$. By the inductive hypothesis, let $p$ be a $(J_{k})_{k<n}$-Laver-tree such that whenever $s\in p$, $\{x\in X_{n}\;|\;s^{\frown}x\in A\}\in J_{n}^{+}$. Let $p^{\prime}$ be the tree of all $n+1$-element sequences $s$ with $s\upharpoonright n\in p$ and $(s\upharpoonright n,s(n))\in A$. Then $p^{\prime}$ is a $(J_{k})_{k\leq n}$-Laver-tree and $T(p^{\prime})\subseteq A$.
On the other hand, suppose that $A\subseteq X_{0}\times\dots\times X_{n}$ and $p$ is a $(J_{k})_{k\leq n}$-Laver tree such that $T(p)\subseteq A$. Whenever $s\in p$, $|s|=n$, then $\operatorname{osucc}_{p}(s)\in J_{n}^{+}$ by assumption. On the other hand, $p^{\prime}:=\{s\in p\;|\;|s|\leq n\}$ clearly is a $(J_{k})_{k<n}$-Laver tree and $T(p^{\prime})=\{s\in p\;|\;|s|=n\}$. Therefore,
$\{(x_{0},\dots,x_{n-1})\in X_{0}\times\dots\times X_{n-1}\;|\;\{x_{n}\in X_{n}\;|\;(x_{0},\dots,x_{n})\in A\}\in J_{n}^{+}\}\supseteq T(p^{\prime})$
Since $T(p^{\prime})\in(J_{0}\otimes\dots\otimes J_{n-1})^{+}$ by induction, $A\in(J_{0}\otimes\dots\otimes J_{n})^{+}$. ∎
@@
where $\subseteq^{*}$ is defined with regard to $I_{0}\otimes\dots\otimes I_{n-1}$.
@@
$\{t\in\operatorname{osucc}_{q_{2}}^{n}(\emptyset)\;|\;\operatorname{osucc}_{q_{2}}(t)\not\subseteq^{*}\operatorname{osucc}_{q_{i}}(t)\}\in I_{0}\otimes\dots\otimes I_{n-1}.$
@@
$\{t\in\operatorname{osucc}_{p}^{n}(\emptyset)\;|\;\operatorname{osucc}_{p}(t)\in I_{n}\}\subseteq A\in I_{0}\otimes\dots\otimes I_{n-1}.$
@@
$\{r\in\operatorname{osucc}_{p}^{n}(t)\;|\;\operatorname{osucc}_{p}^{n}(t^{\frown}r)\in I_{|t|+n}\}\in I_{|t|}\otimes\dots\otimes I_{|t|+(n-1)}.$
Let $q$ consist of all those $t\in q^{\prime}$ such that every initial segment of $t$ is also in $q^{\prime}$. We claim that $q\in\mathbb{L}$ and $\operatorname{stem}(q)=\emptyset$. By Claim 2 and Claim 3, $\emptyset\in q$ and so $q$ is nonempty. It is clearly closed under initial segments by construction. So we have to show that it splits widely enough. Let $t\in q$. By assumption, $\operatorname{osucc}_{p}(t)\in I_{|t|}^{+}$. Additionally, $\operatorname{osucc}_{q}(t)$ consists of all $x\in\operatorname{osucc}_{p}(t)$ such that $\operatorname{osucc}_{p}(t^{\frown}x)\in I_{|t|+1}^{+}$ and for every $n$,
$\{r\in\operatorname{osucc}_{p}^{n}(t^{\frown}x)\;|\;\operatorname{osucc}_{p}(t^{\frown}x^{\frown}r)\in I_{|t|+1+n}\}\in I_{|t|+1}\otimes\dots\otimes I_{|t|+n}.$
@@
$\{r\in\operatorname{osucc}_{p}^{n}(t^{\frown}x)\;|\;\operatorname{osucc}_{p}(t^{\frown}x^{\frown}r)\in I_{|t|+1+n}\}\in(I_{|t|+1}\otimes\dots\otimes I_{|t|+n})^{+}.$
@@
contradicting the fact that $t\in q$, or more precisely, the second requirement for $n=1$. In the second case, we have that the set
@@
$\{x\in[\kappa_{|t|}]^{<\mu^{+}}\;|\;\{r\in\prod_{i\in[|t|+1,|t|+n]}[\kappa_{i}]^{<\mu^{+}}\;|\;(x,r)\in A\}\in(I_{|t|}\otimes\dots\otimes I_{|t|+n})^{+}\}\in I_{|t|}^{+}.$
In other words, $A\in(I_{|t|}\otimes(I_{|t|+1}\otimes\dots\otimes I_{|t|+n}))^{+}=(I_{|t|}\otimes\dots\otimes I_{|t|+n})^{+}$, where the equality follows from Lemma 4.7. But this clearly contradicts the fact that $t\in q$, or more precisely, the second requirement for $n+1$.
@@
The following is our crucial strengthening of the main crux of our earlier work (see [JL25, Theorem 4.3]) It is a variation of the more well-known $\delta-$approximation property – that was introduced implicitly by Mitchell in [Mit72] and explicitly by Hamkins in [Ham03] – which holds for a forcing $\mathbb{P}$ if $\mathbb{P}$ does not add any new function whose restrictions to all $\delta-$sized ground-model sets are in the ground model themselves. On one hand, the assumptions of the theorem are stronger, since wer require the function to be unbounded. On the other hand, its conclusion is also stronger, since it states that any such function does not even have an initial segment which is covered by a small ground-model set. This directly implies that $\kappa^{*}$ does not become approachable. Crucially, it does so without relying on the assumption that the poset does not add any new “short functions into small sets”. This assumption notably does not hold for our poset, since it can be seen straightforwardly that every regular cardinal between $\mu$ and $\sup_{n}\kappa_{n}$ is singularized to have countable cofinality.
@@
$p\upharpoonright(\operatorname{stem}(p)^{\frown}x_{0})\Vdash\dot{F}(\check{i})\in\check{A}_{x_{0}}.$
@@
$(p_{x},\dot{q}_{x})\Vdash\dot{F}(\check{i}_{x})\in\check{A}_{x}.$
@@
$A_{x}:=\{\alpha\;|\;\exists s\in p_{x}(|s|=n_{x}\wedge(p_{x}\upharpoonright s,\dot{q}_{x})\Vdash\dot{F}(\check{i}_{x})=\check{\alpha})\}.$
@@
$(p_{x}^{\prime},\dot{q}_{x}^{\prime})\Vdash\dot{F}(\check{i}_{x})\notin\check{B}.$
@@
Let $G\colon[p]\to\mu$ be the function mapping a branch $b\in[p]$ to $\sup_{n}(i_{b\upharpoonright n})$. For every $\alpha<\mu$, the set $\{b\in[p]\;|\;G(b)<\alpha\}$ is closed in $[p]$ with regards to the tree topology, so by the Rubin-Shelah Theorem (see e.g. [She17, Chapter XI, Lemma 3.5 (1)]) and the fact that every $I_{n}$ is $\mu^{+}-$complete, we can find $p^{\prime}\leq_{0}p$ such that $G$ is bounded in $p_{0}$, say with value $i$. We claim that $(p^{\prime},\dot{q}^{\prime})$ and $i$ work.
@@ 5. Non-approachable Points of a Single Cofinality
Let $\kappa$ be a supercompact cardinal and $l\colon\kappa\to V_{\kappa}$ a Laver function. Let $\mu<\kappa$ be regular. We define the poset $(\mathbb{A}(\mu,\kappa),\leq,\leq_{0})$ as the limit of an Easton-support Magidor iteration $(\mathbb{P}_{\alpha},\dot{\mathbb{Q}}_{\alpha})_{\alpha<\kappa}$ with the following iterands. For $\alpha\leq\mu$, we let $\dot{\mathbb{Q}}_{\alpha}$ be trivial.
@@
Remark 5.4.
In the context of generalized properness, the condition in (3) is referred to as a totally $(M,\mathbb{P})$-generic condition.
@@
In $V[G]$, $N[G]\prec H^{V}(\Theta^{\prime})[G]=H^{V[G]}(\Theta^{\prime})$ and so $M:=N[G]\cap H^{V[G]}(\Theta)$ is closed under $F$. By standard arguments (see e.g. [Jak25, Lemma 5.3.6]), the $\kappa$-cc of $\mathbb{A}$ implies the following:
Claim 1.
The Mostowski-collapse of $N[G]$ is a map from $N[G]$ to $K[\pi[G\cap N]]$. Moreover, whenever $\tau\in N$:
@@
    If $\tau^{G}\in V$, then $\tau^{G}\in N$. In other words, $N[G]\cap V=N$.
@@
    $\pi_{N[G]}(\tau^{G})=(\pi_{N}(\tau))^{\pi[G\cap N]}$. If $\tau^{G}\in V$, then $\pi_{N[G]}(\tau^{G})=\pi_{N}(\tau^{G})$.
We want to verify that $M$ witnesses the conclusion of the theorem.
Let $\nu:=M\cap\kappa$ (which is an inaccessible cardinal), $\gamma:=\sup(M\cap\delta^{*})$ and $\delta_{-}^{*}:=\operatorname{otp}(M\cap\delta^{*})$. Since $N[G]\cap V=N$, $\nu=N\cap\kappa$, $\gamma=\sup(N\cap\delta^{*})$ and $\delta_{-}^{*}=\operatorname{otp}(N\cap\delta^{*})$. Then, by Lemma 1.12, $\delta_{-}^{*}=\operatorname{cf}(\gamma)$ is the successor of a singular cardinal of countable cofinality and $d$ induces a normal coloring $e$ on $\delta_{-}^{*}$ such that in any outer model, $\gamma$ is $d$-approachable if and only if $\delta_{-}^{*}$ is $e$-approachable. For $n\in\omega$, let $\delta_{n}^{-}:=\operatorname{otp}(M\cap\delta_{n})=\operatorname{otp}(N\cap\delta_{n})$. Then each $\delta_{n}^{-}$ is a regular cardinal and $\delta_{-}^{*}=(\sup_{n}\delta_{n})^{+}$.
We fix the following notation: Given $\xi<\kappa$, let $\mathbb{A}(\xi)$ consist of all $f\in\mathbb{A}(\mu,\kappa)$ such that $\operatorname{dom}(f)\subseteq\xi$, ordered as a suborder of $\mathbb{A}(\mu,\kappa)$. It is clear that whenever $\xi<\kappa$, the map $p\mapsto p\upharpoonright\xi$ is a projection from $\mathbb{A}$ to $\mathbb{A}(\xi)$. So, for any $\xi<\kappa$, we can let $G(\xi)$ be the $\mathbb{A}(\xi)$-generic filter induced by $G$. It is clear that $G(\xi)=G\cap\mathbb{A}(\xi)$ whenever $\xi<\kappa$ and that $\pi[G\cap N]=G(\nu)$.
We now prove that $M$ satisfies (2). Clearly, $\mathbb{A}(\nu)$ is $\nu$-cc since $\nu$ is inaccessible. Moreover, $l(\nu)=\pi_{N}((\dot{\mathbb{P}},(\delta_{n})_{n\in\omega}),\delta^{*})=(\pi_{N}(\dot{\mathbb{P}}),(\delta_{n}^{-})_{n\in\omega},\delta^{*}_{-})$.
Claim 2.
The following holds in $V$:
  1. (1)
    $\delta^{*}_{-}>\nu$ is the successor of a singular cardinal of countable cofinality;
  2. (2)
    $(\delta_{n}^{-})_{n\in\omega}$ converges to the predecessor of $\delta^{*}_{-}$;
@@
Proof.
We prove (1), (2) and (3) simultaneously. Since $\Theta$ is sufficiently large, it holds in $H^{V}(\Theta)$ that
  1. (1)
    $\delta^{*}>\kappa$ is the successor of a singular cardinal of countable cofinality;
  2. (2)
    $(\delta_{n}^{-})_{n\in\omega}$ converges to the predecessor of $\delta^{*}$;
  3. (3)
[…]
Since $\pi_{N}$ is an isomorphism, the statements (1), (2) and (3) thus hold in $\pi_{N}(H^{V}(\Theta))$. By Lemma 1.14, $\pi_{N}(H^{V}(\Theta))=H^{V}(\pi_{N}(\Theta))$. Since $\pi_{N}(\Theta)$ is sufficiently large, this implies that all three statements hold in $V$. ∎
@@
$\mathbb{A}(\nu+3)\cong\mathbb{A}(\nu)*\pi_{N}(\mathbb{P})*\mathbb{L}(\check{\mu},(\check{\delta}_{n}^{-},\dot{I}_{n},\dot{B}_{n})_{n\in\omega})*\dot{\operatorname{Coll}}(\check{\mu},\check{\delta}^{*}).$
We now turn to proving the non-$d$-approachability of $\gamma$ in two stages.
Claim 3.
In $V[G(\nu+3)]$, $\operatorname{cf}(\gamma)=\mu$ and $\gamma$ is not $d$-approachable.
Proof.
@@
In $V[G(\nu)]$, $\operatorname{cf}(\gamma)=\delta_{-}^{*}$ is still a regular cardinal and $e$ is still a normal, subadditive coloring on $\delta_{-}^{*}$, since $\mathbb{A}(\nu)$ is $\nu$-$c$. The same is true in $V[G(\nu+1)]$, since $\pi_{N}(\dot{\mathbb{P}})$ is forced to preserve $\delta_{-}^{*}$ and every $\delta_{n}^{-}$. By Lemma 4.11 and Theorem 4, in $V[G(\nu+3)]$, $\operatorname{cf}(\delta_{-}^{*})=\mu$ and $\delta_{-}^{*}$ is not $e$-approachable. Therefore, $\operatorname{cf}(\gamma)=\mu$ and $\gamma$ is not $d$-approachable in the same model. ∎
@@
We now verify that the same holds in the final extension $V[G]$. This extension can be seen as a forcing extension of $V[G(\nu+3)]$ using a Prikry-type poset $(\mathbb{A}(\kappa\smallsetminus\nu+3),\leq,\leq_{0})$ such that $(\mathbb{A}(\kappa\smallsetminus\nu+3),\leq_{0})$ is $\mu-$closed and has direct $\mu$-decidability (using the same argument as for Lemma 5.2 (2)). Since there is a strictly increasing and cofinal function from $\mu$ to $\gamma$ in $V[G(\nu+1)]$, the cofinalities of $\mu$ and $\gamma$ are equal in any outer model. Since $\mu$ is a regular cardinal in $V[G]$, $\operatorname{cf}(\gamma)=\mu$ in $V[G]$.
Claim 4.
In $V[G]$, $\gamma$ is not $d$-approachable.
Proof.
@@
In $V[G(\nu+3)]$, fix a function $\Delta\colon\mu\to\gamma$ which is increasing and cofinal. By Fact 1.6, $\gamma$ is $d$-approachable in $V[G]$ if and only if there is an unbounded subset of $\operatorname{im}(\Delta)$ on which $d$ is bounded. So assume toward a contradiction that $n\in\omega$ and $\dot{\Gamma}$ is an $(\mathbb{A}(\kappa\smallsetminus\nu+3),\leq)$-name for an increasing function from $\mu$ into $\operatorname{im}(\Delta)$ such that $d$ is bounded on $\operatorname{im}(\dot{\Gamma})$ with value $n$, forced by some $a\in\mathbb{A}(\kappa\smallsetminus\nu+3)$.
We now construct an increasing sequence $(\alpha_{i})_{i<\mu}$ of elements of $\mu$ and a $\leq_{0}$-decreasing sequence $(a_{i})_{i<\mu}$ (with $a_{0}\leq_{0}a$) of elements of $\mathbb{A}(\kappa\smallsetminus\nu+3)$ such that for any $i<\mu$, $a_{i}\Vdash\dot{\Gamma}(\check{i})=\check{\Delta}(\check{\alpha}_{i})$. If the sequences have been constructed for $j<i$, first let $a_{i}^{\prime}$ be a $\leq_{0}$-lower bound of $(a_{j})_{j<i}$ (or $a_{i}^{\prime}:=a$ if $i=0$). Then
$a_{i}^{\prime}\Vdash\exists\alpha<\mu(\dot{\Gamma}(\check{i})=\check{\Delta}(\alpha))$
so by the maximum principle and the direct $\mu$-decidability of $(\mathbb{A}(\kappa\smallsetminus\nu+3),\leq,\leq_{0})$ we can find $\alpha_{i}$ and $a_{i}\leq_{0}a_{i}^{\prime}$ such that
$a_{i}\Vdash\dot{\Gamma}(\check{i})=\check{\Delta}(\check{\alpha}_{i}).$
Since $\dot{\Gamma}$ is forced to be increasing and $\Delta$ is increasing, $\alpha_{i}>\alpha_{j}$ for all $j<i$.
@@
Now we claim that $\{\Delta(\alpha_{i})\;|\;i<\mu\}$ is a cofinal subset of $\gamma$ on which $d$ is bounded with value $n$. First observe that, because the sequence $(\alpha_{i})_{i<\mu}$ is increasing, the set $\{\alpha_{i}\;|\;i<\mu\}$ is cofinal in $\mu$ and so, since $\Delta$ is increasing and cofinal, $\{\Delta(\alpha_{i})\;|\;i<\mu\}$ is cofinal in $\gamma$. Lastly, whenever $j<i<\mu$, we have
@@
so $d(\Delta(\alpha_{j}),\Delta(\alpha_{i}))\leq n$. Thus, $\{\Delta(\alpha_{i})\;|\;i<\mu\}$ witnesses that $\gamma$ is $d$-approachable in $V[G(\nu+3)]$ which contradicts Claim 3. ∎